Solve the triangle. 64° a=12.1 b Y= ,b= .c=0 (Round to the nearest tenth as needed.)
Answer #1

Solutions & using triangle Laud & we have 8 +64 + 1 = 180 8 = 180-64-71 [8 = 45° Now noung sine law sinto - sinky - sim45° b